OpenStudy (anonymous):

Solve by the elimination method. 3x - 5y = 16 2x + 3y = -2

hero (hero):

y = (16-3x)/-5 y = (-2-2x)/3 y = y 3(16-3x) = -5(-2-2x) 48-9x = 10 + 10x 19x = 38 x = 2 y = -2